# Deblur4DGS: 4D Gaussian Splatting from Blurry Monocular Video

### v1

@@ 4.2 Continuous Dynamic Gaussians Estimation
Blur-Aware Variable Canonical Gaussians. Existing 4D reconstruction methods generally select a single canonical Gaussians $\mathbf{C}$ across the entire video, which may produce results with missing details in scenes with large motion. To alleviate the issue, we suggest variable the canonical Gaussians as time progresses. In such case, the $k$-th canonical Gaussians $\mathbf{C}_{k}$ is only used for some nearby timestamps thus reducing the difficulty of modeling large motion. One way to achieve this is to uniformly divide the video into $L$ segments and select $\mathbf{C}_{k}$ for $k$-th segment. Although it improves performance, selecting the one corresponding to the sharper frame is better for blur removal. In particular, we first uniformly divide the video into $L$ segments and calculate the blur level $b_{t}$ of dynamic areas for $t$-th frame following  [2, 63]. It can be written as,
$b_{t}=\sum_{(u,v)\in\mathbf{M}_{t}}(\Delta\mathbf{B}_{t}(u,v)-\overline{\Delta\mathbf{B}_{t}})^{2},\vspace{-2mm}$ (7)
where $\mathbf{M}_{t}$ indicates dynamic areas. $\Delta\mathbf{B}_{t}$ is the image Laplacian and $\overline{\Delta\mathbf{B}_{t}}$ is its mean value. The larger $b_{t}$ is, the sharper the frame is. To make the start and end frame of the segment as sharp as possible, we look for the sharp frame among their surrounding $H$ frames and redefine them as the start and end of the current segment. After that, we select the Gaussians corresponding to the sharpest frame in each segment as its canonical ones. Gaussian Deformation. We follow Shape-of-Motion [79] to deform dynamic Gaussians, where a set of rigid transformation matrices are deployed. Let $\{\mathbf{x}_{c},\mathbf{r}_{c},\mathbf{s},\mathbf{o},\mathbf{c}\}$, $\{\mathbf{x}_{t},\mathbf{r}_{t},\mathbf{s},\mathbf{o},\mathbf{c}\}$, and $\{\mathbf{A}_{t},\mathbf{E}_{t}\}$ denote a Gaussian in $\mathbf{C}_{k}$, the ones in $\mathbf{D}_{t}$, and the corresponding transformation matrix, respectively. It can be written as,
$\mathbf{x}_{t}=\mathbf{A}_{t}\mathbf{x}_{c}+\mathbf{E}_{t},\qquad\mathbf{r}_{t}=\mathbf{A}_{t}\mathbf{r}_{c}.\vspace{-2mm}$ (8)
Kindly refer to Shape-of-Motion [79] for details.
@@
$\small\begin{split}\mathbf{D}_{t,i}=\mathbf{w}_{t,i}\odot\mathbf{D}_{t-1}+(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t},\text{ }{i}\in[1,N/2],\\ \mathbf{D}_{t,i}=(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t}+\mathbf{w}_{t,i}\odot\mathbf{D}_{t+1},\text{ }{i}\in[N/2,N].\vspace{-1mm}\end{split}$ (9)
@@
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\mathbf{w}_{t,1}+\frac{i-1}{N-1}\odot\mathbf{w}_{t,N}.\vspace{-2mm}$ (10)
@@
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\frac{\mathbf{w}_{t}}{2}+\frac{i-1}{N-1}\odot(-\frac{\mathbf{w}_{t}}{2}).\vspace{-1mm}$ (11)
Thus, we set learnable parameters $\mathbf{w}_{t}$ in Deblur4DGS for continuous dynamic Gaussians estimation within the exposure time. The canonical Gaussians, Gaussian deformation modules, and $\mathbf{w}_{t}$ are jointly optimized. The reconstruction loss for dynamic areas is similar to Eq. 6.
Figure 3: Illustration of artifacts in dynamic areas. After optimization with Eq. 4, static areas of $\hat{\mathbf{I}}_{t,i}$ are sharp while dynamic areas exist notable artifacts, as shown in the yellow and red boxes in (b), respectively. The proposed regularization terms $\mathcal{L}_{reg}$ (see Eq. 15) improve the visual effect, as shown in (c).
4.3 Regularization Terms
After optimization with Eq. 4, static areas of $\hat{\mathbf{I}}_{t,i}$ are sharp while dynamic areas can with notable artifacts, as shown in the yellow and red boxes of Fig. 3(b) respectively. The reasons are below. (1) Note that multiple solutions exist for the model to fulfill Eq. 4. The most ideal one is that every $\hat{\mathbf{I}}_{t,i}$ is sharp, and the most trivial one is that every $\hat{\mathbf{I}}_{t,i}$ is as blurry as $\mathbf{B}_{t}$. (2) As static areas are consistent across the entire video, the model tends to learn the underlying sharp representation for inter-frame consistency. In other words, the inter-frame consistency implicitly regularizes model optimization. To further validate this, we conduct an experiment that removes the inter-frame consistency by reducing the number of frames to one. In such a case, the static areas are blurry after optimization with Eq. 4, which supports our confirmation. (3) Compared to static areas, the inter-frame consistency in dynamic ones is weaker due to object motion. It may provide insufficient regularization to guide sharp representation learning, thus leading to results with artifacts. To avoid this, we introduce some regularization terms, including exposure regularization $\mathcal{L}_{e}$, multi-frame consistency regularization $\mathcal{L}_{mfc}$, and multi-resolution consistency regularization $\mathcal{L}_{mrc}$.
@@ 4.3 Regularization Terms
$\mathcal{L}_{e}=\texttt{max}(0,\mathbf{\epsilon}-\mathbf{w}_{t}).\vspace{-1mm}$ (12)
@@
$\begin{split}\mathcal{L}_{mfc}=\frac{1}{N-1}&\sum_{i=2}^{N}(\left\|\mathbf{M}_{t,i}\odot(\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}-\hat{\mathbf{I}}_{t,i})\right\|_{1}\\ &\;\;\>+\left\|\mathbf{M}_{t,1}\odot(\hat{\mathbf{I}}_{t,{i\rightarrow 1}}-\hat{\mathbf{I}}_{t,1})\right\|_{1}).\vspace{-2mm}\end{split}$ (13)
$\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}$ and $\hat{\mathbf{I}}_{t,{i\rightarrow 1}}$ are obtained by aligning $\hat{\mathbf{I}}_{t,{i-1}}$ to $\hat{\mathbf{I}}_{t,i}$ and aligning $\hat{\mathbf{I}}_{t,i}$ to $\hat{\mathbf{I}}_{t,1}$ with a pre-trained optical flow network [71], respectively. $\mathbf{M}_{t,i}$ and $\mathbf{M}_{t,1}$ are masks that indicate dynamic areas in $\hat{\mathbf{I}}_{t,i}$ and $\hat{\mathbf{I}}_{t,1}$, respectively.
Third, the blur in the lower resolution is lower level and is easier to remove [28, 75], thus the artifacts are less in models trained with down-sampled blurry video. Taking advantage of this, we impose $\mathcal{L}_{mrc}$ to assist the optimization of high-resolution models with results from low-resolution models. It can be written as,
$\mathcal{L}_{mrc}=||(\mathbf{M}_{t,i})_{\downarrow}\odot((\hat{\mathbf{I}}_{t,i})_{\downarrow}-\texttt{sg}(\hat{\mathbf{I}}_{t,i}^{l}))||_{1}.\vspace{-2mm}$ (14)
@@
[…]
Overall, the proposed regularization terms $\mathcal{L}_{reg}$ can be denoted as,
$\mathcal{L}_{reg}=\lambda_{e}\mathcal{L}_{e}+\lambda_{mfc}\mathcal{L}_{mfc}+\lambda_{mrc}\mathcal{L}_{mrc}.\vspace{-1mm}$ (15)
$\lambda_{e}$, $\lambda_{mfc}$, and $\lambda_{mrc}$ are set to 0.1, 2, 1, respectively. Besides, following Shape-of-Motion [79], we also use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, and the details can be seen in the Suppl.
4.4 Application to Multiple Tasks
@@
5 Experiments
5.1 Experimental Settings
Datasets. We evaluate methods on 6 scenes with significant motion blur from Stereo Blur Dataset [107], where each one contains blurry stereo videos and the corresponding sharp ones. For each scene, we extract 24 frames from the original video and use COLMAP [64] to calibrate the camera extrinsic and intrinsic parameters, following DybluRF [72]. Note that DybluRF [72] conducts experiments on ×$\times 2.5$ down-sampled data. We employ both ×$\times 2.5$ down-sampled ones (i.e., $288\times 512$) and the original ones (i.e., $720\times 1080$).
Training Configurations. We use the blurry image sequences from the left camera for training. For stable optimization, we pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ for $200$ epochs. After that, we jointly optimize the camera motion predictor, $\mathbf{S}$, canonical dynamic Gaussians $\{\mathbf{C}_{k}\}_{k=1}^{K}$, deformable operation $\mathcal{F}$ and exposure time parameters $\{\mathbf{w}_{t}\}_{t=1}^{T}$ for $200$ epochs. The learning rate for the camera motion predictor is set to $5\times 10^{-4}$ and decayed to $1\times 10^{-5}$. The learning rate for $\{\mathbf{w}_{t}\}_{t=1}^{T}$ is set to $1\times 10^{-1}$ and decayed to $1\times 10^{-5}$. The learning rate for $\mathbf{S}$, $\{\mathbf{C}_{k}\}_{k=1}^{K}$ and $\mathcal{F}$ are the same as Shape-of-Motion [79]. $N$ is set to 11. $L$ and $H$ are set to 5 and 3 respectively. $\mathbf{\epsilon}$ is set to $1.0$. All experiments are conducted with PyTorch [57] on one Nvidia GeForce RTX A6000 GPU.
Table 1: Quantitative comparisons for novel view synthesis. $\uparrow$ denotes the higher metric the better, and $\downarrow$ denotes the lower one the better. The best is boldfaced, and the second is underlined.
@@ Methods
DeformableGS [90] 17.82 / 0.722 / 0.292 17.12 / 0.751 / 0.363
4DGaussians [84] 21.94 / 0.833 / 0.180 21.69 / 0.852 / 0.238
E-D3DGS [1] 22.85 / 0.843 / 0.175 22.38 / 0.854 / 0.240
DyBluRF [72] 27.09 / 0.919 / 0.090 26.46 / 0.913 / 0.142
Shape-of-Motion [79] 27.89 / 0.928 / 0.116 27.49 / 0.922 / 0.192
Deblur4DGS (Ours) 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Table 2: Quantitative results that pre-process the blurry video with an image or video deblurring method before 4D reconstruction. ‘None’ denotes no image or video deblurring method being used.
Reconstruction
@@
E-D3DGS [1] None 22.85 / 0.843 / 0.175 22.38 / 0.854 / 0.240
Restormer [92] 23.55 / 0.854 / 0.142 22.77 / 0.867 / 0.186
DSTNet [56] 23.21 / 0.863 / 0.132 22.80 / 0.865 / 0.177
BSSTNet [95] 23.28 / 0.857 / 0.128 23.08 / 0.876 / 0.163
Shape-of-Motion [79] None 27.89 / 0.928 / 0.116 27.49 / 0.922 / 0.192
Restormer [92] 28.39 / 0.941 / 0.080 28.25 / 0.940 / 0.117
DSTNet [56] 28.07 / 0.938 / 0.072 28.22 / 0.941 / 0.104
BSSTNet [95] 28.34 / 0.940 / 0.079 28.24 / 0.943 / 0.100
Deblur4DGS (Ours) None 28.92 / 0.949 / 0.060 28.88 / 0.947 / 0.098
Evaluation Configurations. We evaluate methods on novel-view synthesis, deblurring, frame interpolation, and video stabilization tasks. To evaluate novel-view synthesis and deblurring performance, we use the sharp image sequences of the right and left cameras as ground truth respectively. Due to calibrated camera parameters are not absolutely accurate, the rendering results may be slightly spatially misaligned with the ground truth. We adopt the aligned PSNR [22], SSIM [81] and LPIPS [97] as evaluation metrics, following similar works [3, 12, 85]. Specifically, we align rendering results to the ground truth by a pre-trained optical flow network (i.e., PWC-Net [71]) and then calculate metrics between the aligned results and the ground truth. Simultaneously, we report results that calculate metrics between the original rendering images and the ground truth in the Suppl. As there is no ground truth for frame interpolation and video stabilization tasks, we evaluate method performance with recent no-reference metrics, i.e., CLIPIQA [77] and MUSIQ [26].
Table 3: Quantitative comparisons for deblurring. The first category is supervised image or video deblurring methods, and the second is 4D reconstruction-based ones. The best one in each category is boldfaced, and the second one is underlined.

### v2

@@ Introduction
Blur-Aware Variable Canonical Gaussians. Existing 4D reconstruction methods generally select a single canonical Gaussians $\mathbf{C}$ across the entire video, which may produce results with missing details in scenes with large motion. To alleviate the issue, we suggest varying the canonical Gaussians as time progresses. In such case, the $k$-th canonical Gaussians $\mathbf{C}_{k}$ is only used for some nearby timestamps, thus reducing the difficulty of motion modeling. One way to achieve this is to uniformly divide the video into $L$ segments and select $\mathbf{C}_{k}$ for $k$-th segment. Although it improves performance, selecting the one corresponding to the sharper frame is better for blur removal. In particular, we first uniformly divide the video into $L$ segments and calculate the blur level $b_{t}$ of dynamic areas for $t$-th frame following  (Bansal, Raj, and Choudhury 2016; Ren, Qian, and Chen 2020), i.e.,
$b_{t}=\sum_{(u,v)\in\mathbf{M}_{t}}(\Delta\mathbf{B}_{t}(u,v)-\overline{\Delta\mathbf{B}_{t}})^{2}.$ (7)
$\mathbf{M}_{t}$ indicates dynamic areas. $\Delta\mathbf{B}_{t}$ is the image Laplacian and $\overline{\Delta\mathbf{B}_{t}}$ is its mean value. The larger $b_{t}$ is, the sharper the frame is. To make the start and end frame of the segment as sharp as possible, we look for the sharp frame among their surrounding $H$ frames and redefine them as the start and end of current segment. Finally, we select the Gaussians for the sharpest frame in each segment as its canonical ones.
Gaussian Deformation. We deform dynamic Gaussians with a set of rigid transformation matrices, following Shape-of-Motion (Wang et al. 2024). Let $\{\mathbf{x}_{c},\mathbf{r}_{c},\mathbf{s},\mathbf{o},\mathbf{c}\}$, $\{\mathbf{x}_{t},\mathbf{r}_{t},\mathbf{s},\mathbf{o},\mathbf{c}\}$, and $\{\mathbf{A}_{t},\mathbf{E}_{t}\}$ denote a Gaussian in $\mathbf{C}_{k}$, the ones in $\mathbf{D}_{t}$, and the corresponding transformation matrix, respectively. It can be written as,
$\mathbf{x}_{t}=\mathbf{A}_{t}\mathbf{x}_{c}+\mathbf{E}_{t},\qquad\mathbf{r}_{t}=\mathbf{A}_{t}\mathbf{r}_{c}.$ (8)
@@
$\begin{split}\mathbf{D}_{t,i}=\mathbf{w}_{t,i}\odot\mathbf{D}_{t-1}+(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t},\text{ }{i}\in[1,N/2],\\ \mathbf{D}_{t,i}=(1-\mathbf{w}_{t,i})\odot\mathbf{D}_{t}+\mathbf{w}_{t,i}\odot\mathbf{D}_{t+1},\text{ }{i}\in[N/2,N].\end{split}$ (9)
@@
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\mathbf{w}_{t,1}+\frac{i-1}{N-1}\odot\mathbf{w}_{t,N}.$ (10)
@@
$\mathbf{w}_{t,i}=(1-\frac{i-1}{N-1})\odot\frac{\mathbf{w}_{t}}{2}+\frac{i-1}{N-1}\odot(-\frac{\mathbf{w}_{t}}{2}).$ (11)
Finally, we set learnable parameters $\mathbf{w}_{t}$ for continuous dynamic Gaussians estimation within the exposure time. The canonical Gaussians, Gaussian deformation modules, and $\mathbf{w}_{t}$ are jointly optimized. The reconstruction loss for dynamic areas is similar to eq. 6.
Regularization Terms
After optimization with eq. 4, static areas of $\hat{\mathbf{I}}_{t,i}$ are sharp while dynamic areas can with notable artifacts. The reasons are below. (1) Note that multiple solutions exist for the model to fulfill eq. 4. The most ideal one is that every $\hat{\mathbf{I}}_{t,i}$ is sharp, and the most trivial one is that every $\hat{\mathbf{I}}_{t,i}$ is as blurry as $\mathbf{B}_{t}$. (2) As static areas are consistent across the entire video, the model tends to learn the underlying sharp representation for inter-frame consistency. In other words, the inter-frame consistency implicitly regularizes model optimization. To further validate this, we conduct an experiment that removes the inter-frame consistency by reducing the number of frames to one. In such a case, the static areas are blurry after optimization with eq. 4, which supports our confirmation. (3) Compared to static areas, the inter-frame consistency in dynamic ones is weaker due to object motion. It may provide insufficient regularization to guide sharp representation learning, thus leading to artifacts. To avoid this, we introduce regularization terms $\mathcal{L}_{reg}$, including exposure regularization $\mathcal{L}_{e}$, multi-frame consistency term $\mathcal{L}_{mfc}$, and multi-resolution consistency term $\mathcal{L}_{mrc}$.
@@
$\mathcal{L}_{e}=\texttt{max}(0,\mathbf{\epsilon}-\mathbf{w}_{t}).$ (12)
@@
$\begin{split}\mathcal{L}_{mfc}=\frac{1}{N-1}&\sum_{i=2}^{N}(\left\|\mathbf{M}_{t,i}\odot(\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}-\hat{\mathbf{I}}_{t,i})\right\|_{1}\\ &\;\;\>+\left\|\mathbf{M}_{t,1}\odot(\hat{\mathbf{I}}_{t,{i\rightarrow 1}}-\hat{\mathbf{I}}_{t,1})\right\|_{1}).\end{split}$ (13)
$\hat{\mathbf{I}}_{t,{i-1\rightarrow i}}$ and $\hat{\mathbf{I}}_{t,{i\rightarrow 1}}$ are obtained by aligning $\hat{\mathbf{I}}_{t,{i-1}}$ to $\hat{\mathbf{I}}_{t,i}$ and aligning $\hat{\mathbf{I}}_{t,i}$ to $\hat{\mathbf{I}}_{t,1}$ with a pre-trained optical flow network (Sun et al. 2018), respectively. $\mathbf{M}_{t,i}$ and $\mathbf{M}_{t,1}$ are dynamic masks for $\hat{\mathbf{I}}_{t,i}$ and $\hat{\mathbf{I}}_{t,1}$, respectively.
Third, the blur in the lower resolution is lower level and is easier to remove (Kim, Lee, and Cho 2022; Tao et al. 2018), thus the artifacts are less in models trained with down-sampled blurry video. Taking this advantage, we impose $\mathcal{L}_{mrc}$ to assist the optimization of high-resolution models with results from low-resolution models,i.e.,
[…]
@@
Overall, regularization terms $\mathcal{L}_{reg}$ can be denoted as,
$\mathcal{L}_{reg}=\lambda_{e}\mathcal{L}_{e}+\lambda_{mfc}\mathcal{L}_{mfc}+\lambda_{mrc}\mathcal{L}_{mrc}.$ (15)
$\lambda_{e}$, $\lambda_{mfc}$, and $\lambda_{mrc}$ are set to 0.1, 2, 1, respectively. Besides, following Shape-of-Motion (Wang et al. 2024), we also use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, and the details are in the Sec.B of Suppl.
Application to Multiple Tasks
@@
Table 1: novel view synthesis results on synthetic videos.
@@ Methods
DeformableGS 15.73 / 0.623 / 0.382 15.55 / 0.667 / 0.421
4DGaussians 21.98 / 0.801 / 0.197 21.69 / 0.831 / 0.264
E-D3DGS 23.09 / 0.830 / 0.175 22.46 / 0.844 / 0.258
Shape-of-Motion 26.06 / 0.910 / 0.144 25.81 / 0.897 / 0.246
SplineGS 26.05 / 0.901 / 0.158 24.92 / 0.883 / 0.252
DyBluRF 26.04 / 0.916 / 0.090 25.71 / 0.908 / 0.159
BARD-GS 26.91 / 0.923 / 0.077 26.34 / 0.909 / 0.139
Deblur4DGS (Ours) 27.66 / 0.935 / 0.060 27.16 / 0.927 / 0.123
Table 2: novel view synthesis results on real-world videos.
Methods
[…]
Redmi
[…]
BARD-GS
DeformableGS 0.238 / 25.903 15.63 / 0.781 / 0.361
4DGaussians 0.236 / 26.514 21.32 / 0.863 / 0.221
E-D3DGS 0.257 / 24.967 22.69 / 0.883 / 0.217
Shape-of-Motion 0.277 / 24.538 20.53 / 0.854 / 0.289
SplineGS 0.252 / 32.022 23.93 / 0.899 / 0.197
DyBluRF 0.263 / 34.006 22.71 / 0.878 / 0.192
BARD-GS 0.288 / 35.505 22.69 / 0.874 / 0.177
Deblur4DGS (Ours) 0.356 / 36.756 23.10 / 0.879 / 0.161
[…]
Figure 2: Visual comparisons of novel-view synthesis on real-world videos. Our method produces more photo-realistic details and less visual artifacts in both static and dynamic areas, as marked with yellow and red boxes respectively.
Experiments
Experimental Settings
Training Details. For stable optimization, we pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ for $400$ epochs. After that, we jointly optimize the camera motion predictor, $\mathbf{S}$, canonical dynamic Gaussians $\{\mathbf{C}_{k}\}_{k=1}^{K}$, deformable operation $\mathcal{F}$ and exposure time parameters $\{\mathbf{w}_{t}\}_{t=1}^{T}$ for $200$ epochs. The learning rate for camera motion predictor is set to $5\times 10^{-4}$ and decayed to $1\times 10^{-5}$. The learning rate for $\{\mathbf{w}_{t}\}_{t=1}^{T}$ is set to $1\times 10^{-1}$ and decayed to $1\times 10^{-5}$. The learning rate for $\mathbf{S}$, $\{\mathbf{C}_{k}\}_{k=1}^{K}$ and $\mathcal{F}$ follows Shape-of-Motion (Wang et al. 2024). $N$ is set to 11. $L$ and $H$ are set to 5 and 3 respectively. $\mathbf{\epsilon}$ is set to $1.0$. All experiments are conducted with PyTorch (Paszke et al. 2019) on one Nvidia GeForce RTX A6000 GPU.
Evaluation Configurations. The synthetic data contains 9 scenes with significant motion blur from Stereo Blur Dataset (Zhou et al. 2019), where each scene contains blurry stereo videos and the corresponding sharp ones. Note that DyBluRF (Sun et al. 2024a) conducts experiments on $\times 2.5$ down-sampled data. We evaluate on both $\times 2.5$ down-sampled ones (i.e., $288\times 512$) and the original ones (i.e., $720\times 1080$). For novel-view synthesis and deblurring, the rendering results may be spatially misaligned with ground truth due to the calibration error of camera parameters. Thus, we first freeze the pre-trained 4D model and optimize camera poses by minimizing the photometric error between rendering results and ground truth, and then calculate metrics (i.e., PSNR (Huynh-Thu and Ghanbari 2008), SSIM (Wang et al. 2004) and LPIPS (Zhang et al. 2018b)), following COLMAP-Free 3DGS (Fu et al. 2024). As there is no ground truth for frame interpolation and video stabilization, we employ recent no-reference metrics, i.e., CLIPIQA (Wang, Chan, and Loy 2023) and MUSIQ (Ke et al. 2021). Besides, we evaluate on 6 real-world blurry videos (i.e., Redmi data) captured by a Redmi K50 Ultra smartphone and 12 real-world ones (i.e., BARD-GS data) from BARD-GS (Lu et al. 2025), where each one contains 24 frames. For novel-view synthesis, we employ no-reference metrics (i.e., CLIPIQA and MUSIQ) and full-reference metrics (i.e., PSNR, SSIM and LPIPS) for the two data, respectively. For the other three tasks, we use no-reference metrics (i.e., CLIPIQA and MUSIQ) due to no ground truth.
Comparison with State-of-the-Art Methods
We compare with 7 state-of-the-art methods (i.e., DeformableGS (Yang et al. 2024), 4DGaussians (Wu et al. 2024), E-D3DGS (Bae et al. 2024), Shape-of-Motion (Wang et al. 2024), SplineGS (Park et al. 2025), DyBluRF (Sun et al. 2024a) and BARD-GS (Lu et al. 2025)), where DyBluRF and BARD-GS are designed for 4D reconstruction from blurry monocular video based on NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) respectively.
Table 3: Results that pre-process the blurry video with an image or video deblurring method before 4D reconstruction. ‘None’ denotes no deblurring method being used.
4D
@@
E-D3DGS
None 23.09 / 0.830 / 0.175 22.46 / 0.844 / 0.258
Restormer 23.79 / 0.850 / 0.142 22.87 / 0.864 / 0.198
DSTNet 23.59 / 0.861 / 0.132 22.93 / 0.863 / 0.188
BSSTNet 23.68 / 0.855 / 0.128 23.20 / 0.877 / 0.176
Shape-of-Motion
None 26.06 / 0.910 / 0.144 25.81 / 0.897 / 0.246
Restormer 26.80 / 0.917 / 0.085 26.20 / 0.911 / 0.169
DSTNet 26.60 / 0.915 / 0.082 26.08 / 0.914 / 0.140
BSSTNet 26.78 / 0.916 / 0.080 26.06 / 0.916 / 0.125
Deblur4DGS None 27.66 / 0.935 / 0.060 27.16 / 0.927 / 0.123
Novel-view synthesis. table 1 and table 2 summarize the results. First, methods (i.e., DyBluRF and BARD-GS) that perform 4D reconstruction and motion blur modeling jointly yield overall better performance, especially in LPIPS score. Although SplineGS (Park et al. 2025) gets better PSNR and SSIM scores in BARD-GS data, it produces blurry outputs. It is consistent with the finding in BARD-GS (Wang et al. 2024) that PSNR can sometimes yield higher values even when images appear blurrier. Second, benefiting the explicit 3D representation manner, BARD-GS outperforms DyBluRF, being the most competitive method. Third, compared with BARD-GS, our Deblur4DGS performs better due to the introduction of a series of regularization terms to avoid trivial solutions and blur-aware variable canonical Gaussians to better represent dynamic objects. Visual results in fig. 9 shows that Deblur4DGS removes blur more clearly and produces less visual artifacts in both static and dynamic areas. Per-scene results, more visual results and rendering speed comparisons are in the Sec.C of Suppl.
In addition, to further demonstrate the effectiveness of Deblur4DGS, we first pre-process the blurry videos with state-of-the-art image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang, Xie, and Yao 2024)) deblurring methods and then perform 4D reconstruction, as shown in table 3. Compared with reconstruction from blurry videos, the incorporation of deblurring models improves performance. This is because the deblurring models remove some blur, facilitating sharp scene reconstruction. However, as the deblurring methods cannot perceive 3D structure and maintain scene geometric consistency, the reconstruction results are still unsatisfactory. In contrast, Deblur4DGS jointly reconstructs scene geometry and processes motion blur in 3D space, achieving better scene reconstruction results.
Deblurring. Apart from 4D reconstruction-based methods, we compare with some state-of-the-art image and video deblurring ones. The results are in the Sec.C of Suppl. Deblur4DGS obtains better results than 4D reconstruction-based methods and comparable ones to deblurring-specific ones. Compared with the former, Deblur4DGS better reconstructs the scene, thus performing better. Note that the latter ones are trained on large paired data in supervised manner while Deblur4DGS is optimized with the given blurry video in a self-supervised manner. Although the data prior makes them perform better, Deblur4DGS is more convenient to use.
Frame Interpolation. We interpolate camera poses and timestamps to generate $\times 16$ frame interpolation results. We compare with 4D reconstruction-based methods and some video frame interpolation ones (i.e., RIFE (Huang et al. 2022), EMAVFI (Zhang et al. 2023a), and VIDUE (Shang et al. 2023)). The results are in the Sec.C of Suppl. VIDUE is trained with large paired data in a supervised manner for joint deblurring and fame interpolation, thus achieving better results. Deform4DGS performs satisfactorily overall.
Video Stabilization. We employ a Gaussian filter to smooth camera poses for video stabilization, following  (Peng et al. 2024b). The results are in the Sec.C of Suppl. Deblur4DGS achieves pleasant scores compared with 2D video stabilization methods (i.e., MeshFlow (Liu et al. 2016) and NNDVS (Zhang et al. 2023b)) and 4D reconstruction-based ones, which benefits from better geometry reconstruction.
Figure 3: Effect of continuous camera pose (ECP) and dynamic Gaussian (EDG) estimation.
Ablation Study
We conduct experiments to validate the effectiveness of each strategy. As our strategies in Deblur4DGS are mainly designed to process dynamic areas, we exclude the pixels of static areas to calculate the metrics in dynamic ones.
Effect of ECP and EDG
ECP and EDG are introduced to process camera motion blur and object motion blur, respectively. Quantitative results and visual comparisons are shown in table 4 and fig. 3, respectively. First, without ECP and EDG, the results are almost as blurry as the input frame, as shown in fig. 3(b). Second, only with ECP, the static areas are sharp but may lead to visual artifacts in dynamic areas, as shown in fig. 3(c). It is because ECP cannot simulate the object movement. Third, we further introduce EDG to simulate that, producing visually pleasant results in both areas, as shown in fig. 3(d).
Effect of Regularization Terms
The effect of exposure regularization $\mathcal{L}_{e}$, multi-frame consistency term $\mathcal{L}_{mfc}$ and multi-resolution consistency term $\mathcal{L}_{mrc}$ are in table 5. Visual results are in the Sec.D of Suppl. Without these regularization terms, noticeable artifacts appear in dynamic regions, resulting in degraded performance. By regularizing the object motion within the exposure time distinguished, $\mathcal{L}_{e}$ improves performance. Besides, $\mathcal{L}_{mfc}$ and $\mathcal{L}_{mrc}$ additionally regularize multi-frame and multi-resolution consistency respectively, helping to alleviate artifacts. Their combinations perform best.
Effect of BAV Canonical Gaussians.
The results are summarized in table 6. First, selecting a single canonical Gaussians across the entire video (i.e., None) leads to poor performance, due to the challenge of modeling large object motion. Second, selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates this, leading to performance gain. We also experiment with an optical flow-based strategy (Shaw et al. ) to select canonical Gaussians, performs similar to the uniform selection. This may be due to the inaccurate estimation of optical flow from blurry images. Third, our blur-aware selection is better, as the canonical Gaussians corresponding to the sharper frame help blur removal. Visual results are in the Sec.D of Suppl.
Table 4: Effect about the estimation of continuous camera poses (ECP) and dynamic Gaussians (EDG).
ECP EDG
[…]
$288\times 512$
[…]
$720\times 1080$
$\times$ $\times$ 22.10 / 0.988 / 0.018 22.34 / 0.988 / 0.016
$\checkmark$ $\times$ 22.30 / 0.989 / 0.016 22.39 / 0.988 / 0.015
$\checkmark$ $\checkmark$ 22.36 / 0.989 / 0.015 22.63 / 0.990 / 0.014
Table 5: Effetct of regularization terms (see eq. 15).
$\mathcal{L}_{e}$ $\mathcal{L}_{mfc}$ $\mathcal{L}_{mrc}$
[…]
$288\times 512$
[…]
$720\times 1080$
$\times$ $\times$ $\times$ 21.87 / 0.987 / 0.017 22.30 / 0.988 / 0.016
$\times$ $\checkmark$ $\checkmark$ 22.30 / 0.989 / 0.015 22.56 / 0.989 / 0.015
$\checkmark$ $\times$ $\times$ 22.01 / 0.988 / 0.015 22.40 / 0.989 / 0.015
$\checkmark$ $\checkmark$ $\times$ 22.16 / 0.989 / 0.016 22.49 / 0.989 / 0.015
$\checkmark$ $\times$ $\checkmark$ 22.22 / 0.989 / 0.015 22.54 / 0.989 / 0.014
$\checkmark$ $\checkmark$ $\checkmark$ 22.36 / 0.989 / 0.015 22.63 / 0.990 / 0.014
Table 6: Effect of blur-aware variable (BAV) canonical Gaussians. ‘None’ denotes selecting a single one.